OpenStudy (anonymous):

Given the linear equation y=4/5x-2 , find the y-coordinates of the points (-5, ), (0, ), and (5, ). Please show all of your work. Plot those points and graph the linear equation. Submit your graph through the dropbox.

jimthompson5910 (jim_thompson5910):

I'll do one to get you started When x = -5, we can use this to find y y = 4/5x - 2 y = (4/5)*(-5) - 2 y = (4*(-5))/5 - 2 y = -20/5 - 2 y = -4 - 2 y = -6 So when x = -5, y is y = -6 So we have the point (-5,-6) Do the same with x = 0 and y = 5 to find the other two points.

OpenStudy (anonymous):

so im so confused on this right now

OpenStudy (precal):



OpenStudy (precal):

do you know how to solve equations?

OpenStudy (anonymous):

somewhat but not these i dont understand math im dyslexic in math